Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 590
Entry time: 09 Jan 2019, 10:04
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 09.30 |
|
|
SPB/SFX
- New Karabo release has been installed.
FXE
SQS
SCS
- EET and AE should get in touch to discuss how to speed up the crates fixing. Some details need to be sorted out.
MID
HED
- Mono in XTD6 need some work. Meeting today to discuss details with AE
- HI res mono needs to be opened. Work being organized for next week.
OPT Lasers
- The power glitches l2 nights ago caused a power loss between 1am and 7 am. The seeder in SASE 1 went off. It is being switched on again but this reduces considerably its lifetime. No problems in SASE3. Tha's probably because different protection voltage drops are used (20V/40V) for the fuses. TS will look into that to understand why and whether the situation can be improved.
XRO
- M2 in SASE3 goes back this week and will be pumped during the week end.
- Vac group busy ventin sections to replace imagers,
XPD
VAC
AE
EET
ITDM
DET
CAS
AoB
- Power cut in the SQS/SCS rackroom needed to fix the emergency power off system. 1 hour needed for each rackroom. The redundant power line will not be affected. The following dates have been agreed:
- SCS - MTeichmann will bring the PLCs ina safe state and give confirmation to TSchoen within 1 hour today
- SQS - Power will be cut next Wednesday morning during the chamber switch. Exact time will be agreed next week.
- Optical lasers are not involved in the power cut.
|